Solve the following equations for the unknown.

1)6.2(x - 15.1) = 9.3(-x - 3.2)


2)x/3 - 7/8 = 2/5(x + 15/4)

Answer by checkley77(9423)

1) 6.2(X-15.1)=9.3(-X-3.2)
6.2X-93.62=-9.3X-29.76
6.2X+9.3X=93.62-29.76
15.5X=63.86
X=63.86/15.5
X=4.12 ANSWER.
PROF
6.2(4.12-15.1)=9.3(-4.12-3.2)
6.2*-10.98=9.3*-7.32
-68.076=-68.076